We assume that the relationships in the text below describe an economy. It is a closed economy with a given (fixed) price-level and with a variable interest rate (the interest rate is given with a whole value ex. 10% is 10 and not 0,1).

C = 425 + 0,4 YD

T = 100

G = 140

I = 100 + 0,1 Y – 50r

MD = L(r;Y) = Y – 100r

MS M/P = 200

YD = (Y-T)

  1. Find the equations for IS- and LM curves.
  2. Find the equilibrium income and interest rate
  3. Calculate the private and public saving surplus
  4. Suppose that the government increases public spending to the value 150. Calculate equilibrium for income, interest rate, public, and private savings and explain furthermore the changes in the economy.
Answer #1

LA In a closed econonry, the goods market equilibrium is given by : Y=C+I+G. - putting the respective values of the variablesB) from equation , we have Y = 200 + 100 r. => Y = Y-200 hutting this value of r in equation ③ we have: 13 Y = 1250 - 100 =>Savings, Pi, is given by : Private Pi=Y-C-T. => P, = 725- 675-100 =) Pi = -50 Private Savings is -50 units, ie, Private dissaAY is one th ink 46 - MPC 10 - 10 16.667 AY 1-0.4 ..=).04 = 16,667.... is change in inimnie is by 163667 units (increase in iPenew = T- Gwew = 100-150 5 50 = PzNew = -50... - Budget The new value of public Deficit is 2.50 units. As the governmieut ex